what is the answer to 38.13 divided by 8.2
Answer by Maths68(1474)

38.13 divided by 8.2
=38.13 / 8.2
=(38.13)*(1/8.2)
Remove the decimal
=(3813/100)*(10/82)
=(3813/10)*(1/82)
=3813/820
=4.65